while passing a slower car on the highway, you accelerate from 19.8 m/s to 27.0 m/s in a time of 10.9s. how far do you travel during this time? what is your magnitude of acceleration?
1
Expert's answer
2019-09-20T09:18:16-0400

Solution:

Let's find the acceleration:

"a=\frac{\upsilon_2-\upsilon_1}{t}=\frac{27.0-19.8}{10.9}=0.66" m/s2.

Let's find the distance:

"S=\upsilon_1t+\frac{at^2}{2}=19.8\cdot{10.9}+\frac{0.66\cdot{(10.9)^2}}{2}="

"=215.8+39.2=255" m.

Answer:

S = 255 m, a = 0.66 m/s2.